OpenStudy (anonymous):

Angie was working on solving the exponential equation 23^x = 6; however, she is not quite sure where to start. Using complete sentences, describe to Angie how to solve this equation and how solving would be different if the bases were equal. plz help mee!! @mathstudent55

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

wats ur choices?

OpenStudy (mathstudent55):

Is the equation \(23^x = 6\) ?

OpenStudy (anonymous):

23x=6 Log both sides we dont use log base we use log base 23 because the inverse property of log base 23 of 23 raised to the power of x just equals x log2323x=log236 x=log6/log23

OpenStudy (anonymous):

do u have it?

OpenStudy (anonymous):

so thats the answer? @Albany_Goon

OpenStudy (anonymous):

@Directrix help me plzz!!!

Directrix (directrix):

The question has been answered, hasn't it?

OpenStudy (anonymous):

yes srry i stepped out

OpenStudy (anonymous):

np thanks for your help :) @Albany_Goon

OpenStudy (mathstudent55):

\(23^x = 6\) \(\log 23^x = \log 6\) \(x \log 23 = \log 6\) \(x = \dfrac{\log 6}{\log 23} \)